OpenStudy (anonymous):

how do u solve 1 2/8 - 1/9 =

OpenStudy (anonymous):

Change them so they have a common denominator.. i.e. 1 18/72-8/72 then you subtract the numerator and... voila. your answer is 1 10/72 (you simplify!!)

OpenStudy (anonymous):

There are many ways to solve this problem. Here is what I would do . . . I. First, convert 1 2/8 into an improper fraction: 1 2/8 = (1)(8) + 2 = 10/8 II. Then find a common denominator. In this particular case, the easiest thing to do is the following . . . 10(9)/8(9) - 1(8)/9(8) 90/72 - 8/72 = 82/72 = 1 10/72